OpenStudy (samsan9):

find the vertices and the foci of (x-5)^2/4+(y+3)/16=1.

OpenStudy (samsan9):

\[\frac{ (x-5)^2 }{ 4 }+\frac{ (y+3)^2 }{16 }=1\]

OpenStudy (samsan9):

this?

OpenStudy (anonymous):

Yup one sec. Let me solve it.

OpenStudy (anonymous):

So the first thing you have to do is find the center right? Which is (h,k) so what is the center?

OpenStudy (samsan9):

(5,-3)

OpenStudy (anonymous):

Ok so is it a vertical elipse or horizontal?

OpenStudy (samsan9):

vertical i think?

OpenStudy (anonymous):

Yup so in order to find how high it goes above the center you need to find the square root of 16 which is 4. So for the vertices they will be 4 above and 4 below the center point right?

OpenStudy (samsan9):

yes i think so

OpenStudy (anonymous):

Ok so what are your vertices then?

OpenStudy (samsan9):

(2,4)

OpenStudy (anonymous):

No you have two sets of vertices and so one will be 4 above (5,-3) and one will be below (5,-3) so they would be (5,1) and (5,-7)

OpenStudy (anonymous):

So now you have to find the foci which is represented by the equation c^2=a^2-b^2 and since 16 is A because it's the bigger number and 4 is b because it is the smaller number you just plug them into the equation.

OpenStudy (anonymous):

So what is 16-4?

OpenStudy (samsan9):

12

OpenStudy (anonymous):

Ok and since it's c squared you have to find the square root of 12 but since there is no perfect root c=plus or minus the square root of 12. And that is your foci. So just remember the equation c^2=a^2-b^2 and also that in the equation the bigger number is always a^2 and the smaller number is always b^2.

OpenStudy (anonymous):

So Vertices=(5,1) and (5,-7) Foci=

OpenStudy (anonymous):

plus or minus sqrt of 12

OpenStudy (samsan9):

do you think i could tag you in another problem of mine?

OpenStudy (anonymous):

Sure i'll try to help.

OpenStudy (samsan9):

it might be a little long... but it is one of my last problems on here

OpenStudy (anonymous):

Ok